# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ III-B Asymptotic Performance Analysis
In this subsection, we focus on the performance analysis for the single-user system, including the SNR’s closed-form expression, the lower/upper bounds in the ULA and UPA cases, respectively, and their corresponding asymptotic gains as the antenna number $N$ goes to infinity. For ease of exposition, we assume that the user is located along the $z$-axis (i.e., $\psi=\frac{\pi}{2}$ and $\phi=0$) and its position is denoted by $\mathbf{q}=[0,0,r]^{T}$. In this case, based on the optimal zenith angle obtained in (18a), the entire array region can be divided into inner and outer areas by comparing $\operatorname{arccos}(\vec{\mathbf{q}}_{n}^{T}\mathbf{e}_{3})$ and $\theta_{\mathrm{max}}$, as shown in Fig. 4. Specifically, if RA $n$ is located in the inner area, we have $\operatorname{arccos}(\vec{\mathbf{q}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}}$ and $\epsilon_{n}=0$, i.e., RA $n$ can adjust its boresight to perfectly align with the user direction to obtain the maximum antenna directional gain. Conversely, for RA $n$ located in the outer area, we have $\operatorname{arccos}(\vec{\mathbf{q}}_{n}^{T}\mathbf{e}_{3})>\theta_{\mathrm{max}}$ and $\epsilon_{n}=\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}$, i.e., RA $n$ can only serve the user with $\theta_{\mathrm{z},n}=\theta_{\mathrm{max}}$ and the RA boresight is offset from the user direction by an angle $\epsilon_{n}>0$ due to the zenith angle constraint in (4). Based on the above discussion, the projection between the user direction vector and the optimal pointing vector of RA $n$ can be expressed as
@@
$\gamma\simeq\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi d^{2}}\int_{-\frac{N_{y}d}{2}}^{\frac{N_{y}d}{2}}{\int_{-\frac{N_{x}d}{2}}^{\frac{N_{x}d}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{\frac{1}{r^{2}}(x^{2}+y^{2})}\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (21)
@@
$\tilde{\gamma}=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi d}\int_{-\frac{N_{x}d}{2}}^{\frac{N_{x}d}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}|\frac{x}{r}|-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}.$ (22)
It can be observed that the generic form of the SNR given in (22) is still very complicated for further analysis due to the directivity factor $p$ of the directional gain pattern. In the following, we discuss the typical case of $p=\frac{1}{2}$, i.e., the cosine pattern based on the projected aperture.
@@
${\tilde{\gamma}}=\begin{cases}{\frac{2\bar{P}\xi\delta}{\pi}\triangle_{\mathrm{span}}(N_{x}),}&{N_{x}\leq\hat{N}_{x}}\\ {\frac{2\bar{P}\xi\delta}{\pi}\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\triangle_{\mathrm{span}}(N_{x})-\theta_{\mathrm{max}}\right)\right],}&{N_{x}>\hat{N}_{x},}\end{cases}$ (23)
where $\triangle_{\mathrm{span}}(N_{x})\triangleq\operatorname{arctan}\frac{N_{x}\delta}{2}$ denotes the span angle of the user, which is half of the angle formed by the two line segments connecting the user with both ends of the RA array, as illustrated in Fig. 4, and $\hat{N}_{x}\triangleq 2\left\lfloor\frac{\operatorname{tan}\theta_{\mathrm{max}}}{\delta}\right\rfloor+1$ is the maximum number of antennas in the inner area of the array.
@@
Theorem 1 shows that with the applied MRC receive beamforming and the obtained optimal RA deflection angles, the resultant maximum SNR of the proposed ULA-based RA system scales with the antenna number $N_{x}$ according to the span angle function $\triangle_{\mathrm{span}}(N_{x})$. Furthermore, given the antenna element size $S$, the antenna separation $d$, and the propagation distance $r$, the maximum SNR of the ULA-based RA system mainly depends on the allowable deflection angle range for antenna boresight adjustment and the ULA size.
@@
Remark 1: By applying the linear approximation for the arctangent function, i.e., $\operatorname{arctan}(x)\approx\frac{\pi}{4}x,\;-1\leq x\leq 1$ [35], the SNR obtained in the first case of (23) can be approximated by $\tilde{\gamma}\approx\frac{1}{2}N_{x}\bar{P}\xi\delta$ since we have $0\leq\frac{N_{x}\delta}{2}\leq 1$ when $N_{x}\leq\hat{N}_{x}$. Thus, the resultant SNR increases linearly with the number of RAs when $N_{x}\leq\hat{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})\leq\theta_{\mathrm{max}}$. It can be verified that $f(x)\triangleq\operatorname{sin}\left(\operatorname{arctan}x-\theta_{\mathrm{max}}\right)$ is a concave increasing function with respect to $x$ and $\lim_{x\to\infty}f^{{}^{\prime}}(x)=0$, which indicates that when $N_{x}>\hat{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})>\theta_{\mathrm{max}}$, the growth speed of the maximum SNR slows down as the number of RAs further increases until it reaches zero.
@@
For the infinitely large-scale ULA such that $N_{x}\to\infty$, since $\operatorname{arctan}\frac{N_{x}\delta}{2}\to\frac{\pi}{2}$ as $\frac{N_{x}\delta}{2}\to\infty$, the resultant SNR in (23) reduces to
@@
$\lim_{N_{x}\to\infty}\tilde{\gamma}=\frac{2\xi\delta}{\pi}\bar{P}\left(\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\right).$ (24)
@@
It is observed that a higher asymptotic SNR can be achieved for a ULA-based RA system with a larger allowable deflection angle range $[0,\theta_{\mathrm{max}}]$. By letting $\theta_{\mathrm{max}}=0$ in (24), the asymptotic SNR for the conventional fixed-antenna system is given by [^3] In this paper, we assume that the boresight of each antenna is parallel to the $z$-axis in the fixed-antenna system for ease of exposition.
$\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}=\frac{2\xi\delta}{\pi}\bar{P}.$ (25)
@@
Accordingly, the ratio of the asymptotic SNR of the RA system to that of the fixed-antenna system can be expressed as
$\frac{\lim_{N_{x}\to\infty}\tilde{\gamma}}{\lim_{N_{x}\to\infty}\tilde{\gamma}_{\mathrm{fixed}}}=\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\geq 1,$ (26)
@@
where the inequality holds since $f(x)=x+\operatorname{cos}x$ is a monotonically increasing function with respect to $x$ and $f(0)=1$. Therefore, by exploiting the additional spatial DoFs in terms of 3D boresights to improve the array gain, the proposed RA-enabled system with the optimal deflection angles in (18) will outperform the fixed-antenna system, and the performance gain increases with the maximum allowable zenith angle $\theta_{\mathrm{max}}$.
@@
Lemma 1: If we define the closed-form SNR given in (23) as a function with respect to the span angle of the user, i.e., $\tilde{\gamma}(\triangle_{\mathrm{span}}(N_{x}))$, Theorem 1 can be extended to the general case where the user locates around the ULA with $\phi=\frac{\pi}{2}$ and an arbitrary azimuth angle $\psi\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$. Based on the three possible geometric relationships illustrated in Fig. 5, the resultant SNR for a given azimuth angle $\psi$ is expressed as
@@
${\tilde{\gamma}}=\begin{cases}{\frac{1}{2\operatorname{cos}\psi}\left[\tilde{\gamma}(\triangle_{2})-\tilde{\gamma}(\triangle_{1})\right],}&{\text{Case 1:}\;\psi\in\left[-\frac{\pi}{2},-\frac{N_{x}\delta}{2}\right)}\\ {\frac{1}{2\operatorname{cos}\psi}\left[\tilde{\gamma}(\triangle_{1})+\tilde{\gamma}(\triangle_{2})\right],}&{\text{Case 2:}\;\psi\in\left[-\frac{N_{x}\delta}{2},\frac{N_{x}\delta}{2}\right]}\\ {\frac{1}{2\operatorname{cos}\psi}\left[\tilde{\gamma}(\triangle_{1})-\tilde{\gamma}(\triangle_{2})\right],}&{\text{Case 3:}\;\psi\in\left(\frac{N_{x}\delta}{2},\frac{\pi}{2}\right].}\end{cases}$ (27)
@@
In addition, since all geometric relationships shown in Fig. 5 will reduce to the symmetrical case in Fig. 4 when $N_{x}\to\infty$ regardless of $\psi$, the asymptotic SNR is the same as (24).
@@
Lemma 2: For a UPA-based RA system with $\sqrt{(N_{x}d)^{2}+(N_{y}d)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$ and $\theta_{\mathrm{max}}\leq\frac{\pi}{4}$, we have
@@
Lemma 2 shows that when the entire UPA region is located in the inner area, the SNR with MRC beamforming and optimal RA deflection angles increases linearly with the antenna number $N=N_{x}N_{y}$.
For a large RA array where the UPA region exceeds the inner area, it is challenging to obtain a closed-form expression for (21) due to the double integral and the circular boundary between the inner and outer areas. Alternatively, we first derive its lower/upper bounds for drawing useful insights as follows.
Theorem 2: For the UPA-based RA system, defining $R_{\mathrm{lb}}=\frac{1}{2}\mathrm{min}\{N_{x}d,N_{y}d\}$ and $R_{\mathrm{ub}}=\frac{1}{2}\sqrt{(N_{x}d)^{2}+(N_{y}d)^{2}}$ as the radii of the inscribed and circumscribed disks of the rectangular region $N_{x}d\times N_{y}d$ occupied by the UPA, the resultant SNR is lower-/upper-bounded by
@@
$\left.\quad\quad\quad\quad\quad\int_{\operatorname{arctan}\frac{D}{r}-\theta_{\mathrm{max}}}^{\operatorname{arctan}\frac{R}{r}-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon\right],$ (30)
with $D\triangleq\mathrm{min}\{R,r\operatorname{tan}\theta_{\mathrm{max}}\}$ being the radius of the inner area in which the RAs can adjust their boresights to align with the user direction.
@@
Lemma 3: For $p=\frac{1}{2}$, the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ can be expressed in closed form as (31), shown at the top of the next page, where $\theta_{R}\triangleq\operatorname{arctan}\frac{R}{r}$.
@@
By combining Theorem 2 and Lemma 3, the lower/upper bounds for an RA-enabled system can be obtained according to the UPA size. To obtain the lower/upper bounds for the fixed-antenna system, we set $\theta_{\mathrm{max}}=0$ in (III-B) and calculate the integral in a manner similar to Appendix D, yielding
@@
For the case of $\theta_{\mathrm{max}}>0$, the inequality $\mathcal{G}(R,\frac{1}{2},\theta_{\mathrm{max}})>2\bar{P}\xi\left[1-\operatorname{cos}\theta_{R}+\operatorname{sin}\theta_{\mathrm{max}}\left(\operatorname{ln}\frac{1+\operatorname{sin}\theta_{R}}{\operatorname{cos}\theta_{R}}+\operatorname{ln}\frac{1-\operatorname{sin}\theta_{\mathrm{max}}}{1-\operatorname{sin}^{2}\theta_{\mathrm{max}}}\right)\right]>\mathcal{G}(R,\frac{1}{2},0)$ must hold, indicating that the SNR’s lower bound of the RA system is always higher than that of the fixed-antenna system when $R_{\mathrm{lb}}>r\operatorname{tan}\theta_{\mathrm{max}}$. As the UPA size $N_{x}d,N_{y}d\to\infty$, the radii of the inscribed and circumscribed disks of the UPA region $N_{x}d\times N_{y}d$ go to infinity, i.e., $R_{\mathrm{lb}},R_{\mathrm{ub}}\to\infty$. Therefore, the lower/upper bounds given by Theorem 2 approach to the same limit due to the identical form of the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ as $R\to\infty$. Based on the above, we draw a conclusion that for the UPA setting, the RA system also achieves a higher asymptotic SNR than the fixed-antenna system, like the previous ULA case.
IV Multi-User Case under Multi-path Channel
In this section, we consider the general multi-user and multi-path channel setup, i.e., $K>1$ and $Q\geq 0$. Specifically, an AO algorithm and a two-stage algorithm are proposed to solve problem (P1) suboptimally, which offer different tradeoffs between system performance and computational complexity.

### v4

@@ III-B Asymptotic Performance Analysis
In this subsection, we focus on the performance analysis for the single-user system, and derive a closed-form expression for the SNR in the ULA case, lower/upper bounds for the SNR in the UPA case, and the asymptotic gains in both cases as the antenna number $N$ goes to infinity. For ease of exposition, we assume that the user is located along the $z$-axis (i.e., $\psi=\frac{\pi}{2}$ and $\phi=0$) and its position is denoted by $\mathbf{u}=[0,0,r]^{T}$. In this case, based on the optimal pointing vector obtained in (20), the entire array region can be divided into inner and outer areas by comparing $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})$ and $\theta_{\mathrm{max}}$, as shown in Fig. 3. Specifically, if RA $n$ is located in the inner area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}}$ and $\epsilon_{n}=0$, i.e., RA $n$ can adjust its boresight to perfectly align with the user direction to obtain the maximum antenna directional gain. Conversely, for RA $n$ located in the outer area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})>\theta_{\mathrm{max}}$ and $\epsilon_{n}=\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}$, i.e., RA $n$ can only serve the user with $\theta_{\mathrm{z},n}=\theta_{\mathrm{max}}$ and the RA boresight is offset from the user direction by an angle $\epsilon_{n}>0$ due to the zenith angle constraint in (4). Based on the above discussion, the projection between the user direction vector and the optimal pointing vector of RA $n$ can be expressed as
@@
$\gamma\simeq\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta^{2}}\int_{-\frac{N_{y}\Delta}{2}}^{\frac{N_{y}\Delta}{2}}{\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{\frac{1}{r^{2}}(x^{2}+y^{2})}\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (24)
@@
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta}\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(|\frac{x}{r}|\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}.$ (25)
It can be observed that the SNR given in (25) is still very complicated for further analysis since the directivity factor $p$ exists as a power exponent of the cosine function. In the following, we discuss the typical case of $p=\frac{1}{2}$, i.e., the cosine pattern based on the projected aperture.
@@
${\gamma}=\begin{cases}{\frac{2\bar{P}\xi\delta}{\pi}\triangle_{\mathrm{span}}(N_{x}),}&{N_{x}\leq\bar{N}_{x}}\\ {\frac{2\bar{P}\xi\delta}{\pi}\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\triangle_{\mathrm{span}}(N_{x})-\theta_{\mathrm{max}}\right)\right],}&{N_{x}>\bar{N}_{x},}\end{cases}$ (26)
where $\triangle_{\mathrm{span}}(N_{x})\triangleq\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)$ denotes the span angle of the user, which is the angle formed by the two line segments connecting the user to the center and to one end of the RA array, as illustrated in Fig. 3, and $\bar{N}_{x}\triangleq 2\left\lfloor\frac{\operatorname{tan}\theta_{\mathrm{max}}}{\delta}\right\rfloor+1$ is the maximum number of antennas in the inner area of the array.
@@
Theorem 1 shows that with the applied MRC receive beamforming and the optimized RAs’ pointing vectors, the resultant maximum SNR of the proposed ULA-based RA system scales with the antenna number $N_{x}$ according to the span angle function $\triangle_{\mathrm{span}}(N_{x})$. Furthermore, given the antenna size $A$, the antenna separation $\Delta$, and the propagation distance $r$, the maximum SNR of the ULA-based RA system mainly depends on the allowable range for antenna boresight rotation and the ULA size.
@@
Remark 2: By applying the linear approximation for the arctangent function, i.e., $\operatorname{arctan}(x)\approx\frac{\pi}{4}x,\;-1\leq x\leq 1$ [42], the SNR obtained in the first case of (26) can be approximated by $\gamma\approx\frac{1}{2}N_{x}\bar{P}\xi\delta$ since we have $0\leq\frac{N_{x}\delta}{2}\leq 1$ when $N_{x}\leq\bar{N}_{x}$. Thus, the resultant SNR increases linearly with the number of RAs when $N_{x}\leq\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})\leq\theta_{\mathrm{max}}$. It can be verified that $f(x)\triangleq\operatorname{sin}\left(\operatorname{arctan}\left(x\right)-\theta_{\mathrm{max}}\right)$ is a concave increasing function with respect to $x$, and that $\lim_{x\to\infty}f^{{}^{\prime}}(x)=0$. This indicates that when $N_{x}>\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})>\theta_{\mathrm{max}}$, the growth rate of the maximum SNR gradually decreases as the number of RAs further increases, eventually approaching zero.
@@
For the infinitely large-scale ULA such that $N_{x}\to\infty$, since $\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)\to\frac{\pi}{2}$ as $\frac{N_{x}\delta}{2}\to\infty$, the resultant SNR in (26) reduces to
@@
$\lim_{N_{x}\to\infty}\gamma=\frac{2\xi\delta}{\pi}\bar{P}\left(\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\right).$ (27)
@@
It is observed that a higher asymptotic SNR can be achieved for a ULA-based RA system with a larger rotational range. By letting $\theta_{\mathrm{max}}=0$ in (27), the asymptotic SNR for the conventional fixed-antenna system is given by
$\lim_{N_{x}\to\infty}\gamma_{\mathrm{fixed}}=\frac{2\xi\delta}{\pi}\bar{P},$ (28)
@@
where the boresight of each antenna is assumed to be parallel to the $z$-axis for ease of exposition. Accordingly, the ratio of the asymptotic SNR of the RA system to that of the fixed-antenna system can be expressed as
$\frac{\lim_{N_{x}\to\infty}\gamma}{\lim_{N_{x}\to\infty}\gamma_{\mathrm{fixed}}}=\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\geq 1,$ (29)
@@
where the inequality holds since $f(x)=x+\operatorname{cos}x$ is a monotonically increasing function with respect to $x$ and $f(0)=1$. Therefore, by exploiting the additional spatial DoFs in terms of boresight rotation to improve the array gain, the proposed RA system with the optimal pointing vectors in (20) will outperform the fixed-antenna system, and the performance gap increases with the allowable range of RA’s boresight rotation, as constrained by (4).
@@
Lemma 1: If we define the closed-form SNR given in (26) as a function with respect to the span angle of the user, i.e., $\gamma(\triangle_{\mathrm{span}}(N_{x}))$, Theorem 1 can be extended to the general case where the user is located around the ULA with $\psi=\frac{\pi}{2}$ and an arbitrary azimuth angle $\phi\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$. Based on the three possible geometric relationships illustrated in Fig. 4, the resultant SNR for a given azimuth angle $\phi$ is expressed as
@@
${\tilde{\gamma}}=\begin{cases}{\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{2})-\gamma(\triangle_{1})\right],}&{\text{Case 1:}\;\phi\in\left[-\frac{\pi}{2},-\frac{N_{x}\delta}{2}\right)}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})+\gamma(\triangle_{2})\right],}&{\text{Case 2:}\;\phi\in\left[-\frac{N_{x}\delta}{2},\frac{N_{x}\delta}{2}\right]}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})-\gamma(\triangle_{2})\right],}&{\text{Case 3:}\;\phi\in\left(\frac{N_{x}\delta}{2},\frac{\pi}{2}\right].}\end{cases}$ (30)
@@
In addition, since all geometric relationships shown in Fig. 4 will reduce to the symmetrical case in Fig. 3 when $N_{x}\to\infty$ regardless of $\phi$, the asymptotic SNR turns out to be the same as (27).
@@
Lemma 2: For a UPA-based RA system with $\sqrt{(N_{x}\Delta)^{2}+(N_{y}\Delta)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$ and $\theta_{\mathrm{max}}\leq\frac{\pi}{4}$, we have
@@
Lemma 2 shows that when the entire UPA region is located in the inner area, the optimal SNR in (31) increases linearly with the antenna number $N=N_{x}N_{y}$.
For a large RA array where the UPA region exceeds the inner area, it is challenging to obtain a closed-form expression for (24) due to the double integral and the circular boundary between the inner and outer areas. Alternatively, we first derive its lower/upper-bounds for drawing useful insights as follows.
Theorem 2: For the UPA-based RA system, defining $R_{\mathrm{lb}}=\frac{1}{2}\mathrm{min}\{N_{x}\Delta,N_{y}\Delta\}$ and $R_{\mathrm{ub}}=\frac{1}{2}\sqrt{(N_{x}\Delta)^{2}+(N_{y}\Delta)^{2}}$ as the radii of the inscribed and circumscribed disks of the rectangular region $N_{x}\Delta\times N_{y}\Delta$ occupied by the UPA, the resultant SNR is lower/upper-bounded by
@@
$\left.\quad\quad\quad\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon\right],$ (33)
with $D\triangleq\mathrm{min}\{R,r\operatorname{tan}\theta_{\mathrm{max}}\}$ being the radius of the inner area within which the RAs can adjust their boresights to precisely align with the user direction.
@@
Lemma 3: For $p=\frac{1}{2}$, the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ can be expressed in closed form as (34), shown at the top of the next page, where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$.
@@
By combining Theorem 2 and Lemma 3, the lower/upper bounds for an RA system can be obtained according to the UPA size. To obtain the lower/upper bounds for the fixed-antenna system, we set $\theta_{\mathrm{max}}=0$ in (III-B) and calculate the integral in a manner similar to Appendix D, yielding
@@
For the case of $\theta_{\mathrm{max}}>0$, the inequality $\mathcal{G}(R,\frac{1}{2},\theta_{\mathrm{max}})>2\bar{P}\xi\left[1-\operatorname{cos}\theta_{R}+\operatorname{sin}\theta_{\mathrm{max}}\left(\operatorname{ln}\frac{1+\operatorname{sin}\theta_{R}}{\operatorname{cos}\theta_{R}}+\operatorname{ln}\frac{1-\operatorname{sin}\theta_{\mathrm{max}}}{1-\operatorname{sin}^{2}\theta_{\mathrm{max}}}\right)\right]>\mathcal{G}(R,\frac{1}{2},0)$ always holds, indicating that the SNR’s lower bound of the RA system is always higher than that of the fixed-antenna system when $R_{\mathrm{lb}}>r\operatorname{tan}\theta_{\mathrm{max}}$. As the UPA size $N_{x}\Delta,N_{y}\Delta\to\infty$, the radii of the inscribed and circumscribed disks of the UPA region $N_{x}\Delta\times N_{y}\Delta$ go to infinity, i.e., $R_{\mathrm{lb}},R_{\mathrm{ub}}\to\infty$. Therefore, the lower/upper bounds given by Theorem 2 approach to the same limit due to the identical form of the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ as $R\to\infty$. Based on the above, we draw a conclusion that for the UPA setting, the RA system also achieves a higher asymptotic SNR than the fixed-antenna system, as in the previous ULA case.
IV Multi-User Case under Multipath Channel
In this section, we consider the general multi-user and multipath channel setup, i.e., $K>1$ and $Q\geq 0$. Specifically, an AO algorithm and a two-stage algorithm are proposed to solve problem (P1) suboptimally, which offer different trade-offs between system performance and computational complexity.